# NEAT: Neighborhood-Guided, Efficient, Autoregressive Set Transformer for 3D Molecular Generation

### v1

@@ 3 Methods
with random time steps $\mathbf{t}\in\mathcal{U}(0,1)^{|V_{T}|}$, yielding a set of interpolated positions $\mathbf{X_{t}}$ with velocities $\mathbf{V_{t}}=\mathbf{X}_{1}-\mathbf{X}_{0}$. We then train the conditional vector field $\mathbf{V}_{\theta}$ through regression against the set of CFM path velocities:
$\mathcal{L_{\text{CFM}}}=\frac{1}{|V_{T}|}\lVert\mathbf{V}_{\theta}(\mathbf{X_{t}},\textbf{t},\mathbf{a}[V_{T}],\mathbf{z}_{n})-\mathbf{V_{t}}\rVert_{2}^{2}$ (5)
Note that this equation does not require sequential information as input. The vector field $\mathbf{V}_{\theta}$ is parametrized with the adaptive layer MLP from (Li2024AutoregressiveImageGenerationWithoutVQ).
@@
In order to reduce the transport cost between the source distribution $p_{0}$ and the target distribution $p_{1}$, we employ a simple coupling scheme. Given the positions $\mathbf{X}_{1}$ and $\textbf{X}_{0}$, we solve the linear assignment problem on the distance matrix $\textbf{X}_{1}^{T}\textbf{X}_{0}$ and permute the positions in $\textbf{X}_{0}$ accordingly. This procedure avoids crossing paths, resulting in improved performance.
@@
The molecular generation process begins by randomly sampling an initial atom type $a_{1}$ from the atom type vocabulary $\mathcal{A}$ and drawing its initial position $\mathbf{x}_{1}\in\mathbb{R}^{3}$ from a normal distribution $\mathcal{N}(0,\sigma^{2}I)$. At each subsequent autoregressive step, the current atom set $(\mathbf{a}_{1:n},\mathbf{X}_{1:n})$ is passed through the transformer, producing the set-level representation $\mathbf{z}_{n}$. The next atom type $a_{n+1}$ is predicted from $\mathbf{z}_{n}$ using a linear layer followed by the softmax function. A random initial position $\mathbf{x}_{n+1}$ is then sampled from $\mathcal{N}(0,\sigma^{2}I)$, and the flow ODE is integrated from $t=0$ to $t=1$ using Euler integration to compute the next atom position. This iterative process continues until a stop token is predicted, signaling the end of the molecular generation. The workflow is illustrated in Figure 2. Molecular generation is efficiently implemented using batched operations, enabling parallel generation and high computational speed. Although bidirectional attention precludes the use of query-value caching, our generation process remains fast in practice.
Figure 2: Overview of the proposed model inference workflow.
@@ 4 Experiments
We trained NEAT on the QM9 dataset, a collection of small organic molecules composed of hydrogen (H), carbon (C), nitrogen (N), oxygen (O), and fluorine (F) atoms. The dataset includes molecules with up to 9 heavy atoms (C, N, O, F) and a total of 29 atoms when hydrogen is included (Ramakrishnan2014QM9). Following Hoogeboom2022EDM, we split the dataset into 100,000 molecules for training, 17,748 for validation, and 13,083 for testing.
@@
To evaluate the quality of generated molecules, we adhere to standard practices (Hoogeboom2022EDM; Zhang2025SymDiff; Cheng2025Quetzal) and report atomic and molecular stability, molecular validity, and molecular uniqueness. An atom is considered stable if its valence is valid (i.e., it satisfies the octet or duet rule given its formal charge and environment). A molecule is defined as stable if all of its atoms are stable. Molecular validity is assessed using two approaches: (1) the widespread EDM pipeline (Hoogeboom2022EDM) based on bond lookup tables; (2) the pipeline introduced by Daigavane2024Symphony and adopted by Cheng2025Quetzal, where a molecule is deemed valid if bonds can be assigned using RDKit’s MolFromXYZBlock function and rdDetermineBonds module (version 2023.3.3), which uses the xyz2mol package based on the work of Kim2015xyz2mol. Molecular uniqueness is the percentage of distinct canonical SMILES strings derived from valid molecules and can therefore be thought of as the percentage of valid and unique designs. For all metrics, we generate 10,000 samples and report the mean value and 95% confidence interval across three runs.
@@
We compare NEAT to three autoregressive 3D molecular generation models (G-SchNet (Gebauer2019GSchnNet), Symphony (Daigavane2024Symphony), QUETZAL (Cheng2025Quetzal)) and one model-agnostic method to convert 3D molecular structures into SE(3)-invariant 1D discrete sequences that is used in combination with language models to generate 3D molecules (Geo2Seq (Li2025Geo2Seq)). Results can be found in Table 1. Additional results comparing NEAT to diffusion models can be found in the Appendix. To ensure fair comparisons, we utilized the source code and pre-trained models of the baseline methods, where available, to generate molecules and evaluate their performance using our pipeline. Random examples of generated molecules are provided in Figure 3.
Table 1: Performance comparison of NEAT with autoregressive models trained on the QM9 dataset. All values are given in percent. The best and second-best performing models are highlighted. Models valuated using our pipeline are marked with a star (*). Metrics are reported as the mean and 95% confidence interval over three runs, each generating 10,000 molecular designs. Results in the top row are computed on the QM9 test set of 13,083 molecules.
metrics
atom
stab.
mol.
stab.
lookup
valid
lookup
unique
xyz2mol
valid
xyz2mol
unique
QM9 99.4 95.3 97.7 97.7 100 100
G-SchNet 95.7 68.1 85.5 80.3 75.0 72.5
Symphony 90.8 43.9 68.1 66.5 83.5 81.8
Geo2Seq 98.9 93.2 97.1 81.7 n.a. n.a.
QUETZAL* 98.3 87.5 94.6 90.1 99.3 94.4
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.1 $\pm$ 0.1
NEAT* (ours) 98.2 84.7 93.4 88.6 98.5 94.8
$\pm$ 0.1 $\pm$ 0.4 $\pm$ 0.3 $\pm$ 0.4 $\pm$ 0.1 $\pm$ 0.1
@@
Figure 3: Randomly selected examples of generated molecules (white: hydrogen, gray: carbon, purple: nitrogen, red: oxygen).
Model performance
Model rankings depend on how bonds are inferred from generated point clouds. We report metrics using the lookup method proposed by EDM because it is widely adopted and thus facilitates fair comparisons. A notable limitation is the absence of aromatic bond handling and high sensitivity to bond lengths (Daigavane2024Symphony), which yields systematically lower scores for valid aromatic structures and thus reduces the informative value of these metrics. We further report molecular validity and uniqueness using the xyz2mol method. The conceptually related G-SchNet and Symphony models underperform in terms of all metrics. Geo2Seq obtains the highest scores for atom stability, molecular stability, and (lookup) molecular validity, followed by QUETZAL. However, Geo2Seq performs poorly on molecular uniqueness, which is the metric we consider most critical for downstream applications such as drug discovery. Valid designs are of limited value if they do not span a broad chemical space, and low molecular uniqueness might indicates overfitting to the training data. NEAT outperforms all models in terms of xyz2mol uniqueness, and is a close second to QUETZAL on lookup uniqueness and xyz2mol validity. Additionally, unlike QUETZAL and Geo2Seq, NEAT enforces atom-level permutation invariance and avoids sequential positional encodings, which is essential for prefix completion in molecular design. NEAT’s runtime for generating 10,000 molecules is 105 seconds, slightly faster than QUETZAL’s 134 seconds, indicating that bidirectional attention introduces no practical overhead relative to causal attention. Runtimes are averaged over three runs on a system with an AMD Ryzen 9 7950X 16-core CPU and an NVIDIA GeForce RTX 4090 GPU.
@@
Our model inherits standard limitations of autoregressive decoders. Error propagation might degrade sampling for larger molecules, which could be mitigated by a self-correction mechanism in future work. Our model produces variable-sized molecules, but does not extrapolate beyond the maximum atom count seen during training. For large molecules, pooling node embeddings may cause oversquashing and lose key information. Because bonds are not predicted explicitly, bond assignments still rely on predefined rules. Finally, despite random rotations during training, the lack of $SE(3)$ equivariance can yield orientation-dependent predictions.
@@ 5 Conclusion
We introduce a new method for training an autoregressive transformer with bidirectional attention for 3D molecular generation using neighborhood guidance. Instead of sequential positional encodings, we rely on a subgraph sampling schema for model supervision. Treating the input as a permutation-invariant set removes the need for sequence canonicalization and makes the model well-suited for prefix completion. On QM9, our approach achieves competitive performance, while offering permutation-invariance with respect to atoms and fast sampling. We believe this provides a strong foundation for future work in conditional modeling and autoregressive models.
@@
Code and model weights will be made publicly available upon acceptance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ 3 Methods
with random time steps $\mathbf{t}\in\mathcal{U}(0,1)^{|V_{T}|}$, yielding a set of interpolated positions $\mathbf{X_{t}}$ with velocities $\mathbf{V_{t}}=\mathbf{X}_{1}-\mathbf{X}_{0}$. We then train the conditional vector field $\mathbf{V}_{\theta}$ through atom-wise regression against the set of CFM path velocities:
$\mathcal{L_{\text{CFM}}}=\mathbb{E}_{\mathbf{t},\mathbf{X}_{0},\mathbf{X}_{1}}\left[\frac{1}{|V_{T}|}\lVert\mathbf{V}_{\theta}(\mathbf{X_{t}},\textbf{t},\mathbf{a}[V_{T}],\mathbf{z}_{n})-\mathbf{V_{t}}\rVert_{F}^{2}\right]$ (5)
where $\lVert\cdot\rVert_{F}^{2}$ denotes the squared Frobenius norm. The vector field $\mathbf{V}_{\theta}$ is parametrized with the diffusion transformer (DiT)-style adaptive layer norm (adaLN) MLP from (Li et al., 2024). Input to the MLP for velocity regression is the atom-wise summation of positional encoding, atom type embedding, time step embedding, and the respective source set representation. The final loss is obtained by adding the cross-entropy loss and the flow matching loss.
@@
In order to reduce the transport cost between the source distribution $p_{0}$ and the target distribution $p_{1}$, we employ a simple coupling scheme. Given the positions $\mathbf{X}_{1}$ and $\textbf{X}_{0}$, we solve the linear assignment problem on the Euclidean distance matrix $D(\textbf{X}_{1},\textbf{X}_{0})$ and permute the positions in $\textbf{X}_{0}$ accordingly. This procedure avoids crossing paths, resulting in improved performance.
@@
The generation process begins by randomly sampling an initial atom type $a_{1}$ from the atom type vocabulary $\mathcal{A}$ and drawing its initial position $\mathbf{x}_{1}\in\mathbb{R}^{3}$ from a normal distribution $\mathcal{N}(0,\sigma^{2}I)$. At each subsequent autoregressive step, the current atom set $(\mathbf{a}_{1:n},\mathbf{X}_{1:n})$ is passed through the transformer, producing the set-level representation $\mathbf{z}_{n}$. Note that the positions are not re-centered between autoregressive steps. The next atom type $a_{n+1}$ is predicted from $\mathbf{z}_{n}$ using a linear layer followed by the softmax function. A random initial position $\mathbf{x}_{n+1}$ is then sampled from $\mathcal{N}(0,\sigma^{2}I)$, and the transport equation is integrated from $t=0$ to $t=1$ to compute the next atom position (details about the integration scheme are provided in Appendix A.3). This iterative process continues until a stop token is predicted, signaling the end of the molecular generation. The workflow is illustrated in Figure 2. Molecular generation is efficiently implemented using batched operations, enabling parallel generation and high computational speed.
@@ 4 Experiments
We used the QM9 (Ramakrishnan et al., 2014) and GEOM-Drugs (Axelrod and Gómez-Bombarelli, 2022) datasets, detailed information about the data processing is provided in the Appendix A.2. QM9 is a collection of small organic molecules composed of up to nine heavy atoms (carbon, nitrogen, oxygen, and fluorine), with a total of up to 29 atoms when including hydrogen. We obtained the QM9 dataset from Ramakrishnan et al. (2019) and, following standard practice (Hoogeboom et al., 2022; Cheng et al., 2025), randomly split it into training, validation, and test sets, yielding 100,000 molecules for training, 11,619 for validation, and 12,402 for testing after preprocessing. GEOM-Drugs contains 304,313 drug-like molecules, each with up to 30 conformers. Following Dunn and Koes (2025), we loaded the split data from Koes (2024). In line with Vignac et al. (2023), we retained, for each molecule, the five lowest-energy conformers when available. After preprocessing, the dataset contained 1,172,436 conformers across 243,451 unique molecules for training, 146,394 conformers across 30,430 unique molecules for validation, and 146,788 conformers across 30,432 unique molecules for testing. The mean and median number of conformers per molecule are 4.8 and 5.0, respectively. In the following, we refer to GEOM-Drugs as GEOM.
Model training
We trained on QM9 for 10,000 epochs and on GEOM for 1,000 epochs, keeping the checkpoint with the lowest total validation loss for inference. Training was done on an NVIDIA RTX Pro 6000 Blackwell GPU with batch sizes of 2048 (QM9) and 512 (GEOM). System specifications, implementation details, and hyperparameters of the final models are provided in Appendix A.3. Appendix A.5 presents scaling analyses with respect to transformer and projection head size. Our best-performing GEOM model employs a 12-layer transformer with 12 heads and hidden dimension 768, and an MLP projector comprising 6 AdaLN blocks with hidden dimension 1536 (165M parameters). This configuration offers a favorable trade-off between generation quality and model size. Our best performing QM9 model (282M parameters) uses a 16-layer transformer with 16 heads and hidden dimension 1024. We speculate that having to use a larger model for the seemingly easier QM9 dataset might be due to the highly constrained molecular geometries of the QM9 dataset, which might need higher model capacity.
@@
To evaluate the quality of generated molecules, we adhere to standard practices (Hoogeboom et al., 2022; Daigavane et al., 2024; Zhang et al., 2025; Cheng et al., 2025) and report atomic stability and molecular stability, validity, uniqueness, and novelty. An atom is considered stable if its valence satisfies the octet or duet rule. A molecule is defined as stable if all of its atoms are stable. Molecular uniqueness is the percentage of distinct canonical SMILES strings derived from valid molecules, or in other words, the percentage of valid and unique designs. Novelty is the percentage of valid, unique designs that are not present in the training set. For all metrics, we generated 10,000 samples and reported the mean value and 95% confidence interval across three runs. Random examples of molecules generated by NEAT trained on GEOM are shown in Figure 3. Appendix A.5 provides further examples and a comparison of the distribution of molecular properties between the generated molecules and the training set.
Molecule building
Molecular validity depends on the algorithm used to construct molecules from the 3D clouds of unconnected atoms that are output by most 3D molecular generation models. We assess molecular validity using two approaches: (1) the widespread EDM pipeline (Hoogeboom et al., 2022) based on bond lookup tables, (2) the pipeline introduced by Daigavane et al. (2024) and adopted by Cheng et al. (2025), where a molecule is deemed valid if bonds can be assigned using RDKit’s MolFromXYZBlock function and rdDetermineBonds module, which uses the xyz2mol package based on the work of Kim and Kim (2015). A notable limitation of the EDM pipeline is the absence of aromatic bond handling and a high sensitivity to bond lengths (Daigavane et al., 2024), which yields systematically lower scores for valid aromatic structures and thus reduces the informative value of these metrics. We also observed that the output of RDKit’s rdDetermineBonds module depends on the RDKit version. Throughout this work, we use RDKit v2025.9.1. A detailed discussion of limitations arising from these bond-assignment-based evaluation pipelines is provided in Appendix A.4.
@@
We compare NEAT trained on QM9 to five autoregressive 3D molecular generation models (G-SchNet (Gebauer et al., 2019), G-SphereNet (Luo and Ji, 2022), Symphony (Daigavane et al., 2024), QUETZAL (Cheng et al., 2025), and Geo2Seq (Li et al., 2025b)). The results are reported in Table 1, with additional comparisons to diffusion-based 3D molecular generation models provided in Appendix A.5. We compare NEAT trained on GEOM to QUETZAL and Geo2Seq, since other autoregressive models were not trained on GEOM and report the results in Table 2. Since the evaluation metrics are sensitive to the procedure used to infer bonds from atomic point clouds, including the choice of algorithm, its hyperparameters, and the software’s version, we generated molecules with the source code and model weights of each baseline wherever possible and evaluated them using a unified pipeline. For G-SchNet, G-SphereNet, and Symphony, we used the pre-computed XYZ files released by the authors of Symphony (Daigavane et al., 2023) and computed all metrics with our pipeline. For Geo2Seq, we report the metrics from the original publication.
Unconditional generation performance
On QM9, NEAT outperformed all autoregressive models in terms of xyz2mol molecular validity, uniqueness, and novelty. NEAT ranked second to QUETZAL on lookup uniqueness, and third on lookup validity, atom stability, and molecular stability. While Geo2Seq achieved the highest scores for atom stability, molecular stability, and lookup validity, it performed poorly on molecular uniqueness, which we regard as the most critical metric for downstream applications such as drug discovery. Valid designs are of limited practical value if they do not span a broad chemical space. Moreover, low molecular uniqueness may indicate overfitting to the training distribution. On the GEOM dataset, NEAT achieved the highest lookup molecular validity and uniqueness. Both QUETZAL and NEAT exhibited complete novelty and similarly low molecular stability, close to that of the GEOM test set (2.8%). Further discussion of this value is provided in Appendix A.4. On the xyz2mol metrics, NEAT underperformed relative to QUETZAL, likely due to differences in objective complexity. QUETZAL employs a standard causal transformer objective, whereas NEAT models the full distribution of token continuations, which introduces substantially higher degrees of freedom and increases the risk of error accumulation during inference. Nevertheless, our approach yields greater robustness in downstream applications, a trade-off examined in the subsequent prefix-completion experiments.
Table 1: Performance comparison of NEAT against autoregressive 3D molecular generators trained on QM9. All metrics are reported as percentages. The first row reports results on the QM9 test set (12,402 molecules). All other rows report results computed from 10,000 generated molecules. The best and second-best entries are shown in bold and underlined, respectively. Entries for which generated molecules could be retrieved to be re-evaluated with our pipeline are marked with a single star (*). Entries for which the source code and model weights were available to generate molecules are marked with two stars (**). For these models and for NEAT, we report the mean and 95% confidence interval over three runs.
atom stable mol stable lookup valid lookup unique xyz2mol valid xyz2mol unique xyz2mol novel
QM9 99.3 94.8 97.5 97.5 100.0 100.0 99.8
G-SchNet* 93.7 63.6 80.1 74.7 73.0 70.6 55.1
G-SphereNet* 67.8 14.0 16.4 3.8 37.7 7.2 6.7
Symphony* 90.8 43.9 68.1 66.5 77.0 75.2 41.3
Geo2Seq 98.9 93.2 97.1 81.7 n.a. n.a. n.a.
QUETZAL** 98.3 87.5 94.6 90.1 93.7 89.1 24.7
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.3 $\pm$ 0.2 $\pm$ 0.3
NEAT (ours) 98.1 84.2 93.9 88.8 94.4 90.5 63.9
$\pm$ 0.1 $\pm$ 0.3 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.7
Table 2: Performance of NEAT and other autoregressive 3D molecular generators trained on GEOM on the tasks of unconditional generation and prefix completion. Metrics are reported as percentages. The first row presents metrics computed from the GEOM test set (146,788 conformers of 30,432 unique molecules). All other rows summarize metrics computed from 10,000 generated molecules. For each experiment, the best results are highlighted in bold. Since QUETZAL’s source code and model weights are available, we use them to generate molecules and evaluate them using our pipeline. For QUETZAL and NEAT, we report the mean and 95% confidence interval over three independent runs. For Geo2Seq, metrics are taken directly from the original publication.
atom stable mol stable lookup valid lookup unique xyz2mol valid xyz2mol unique xyz2mol novel
GEOM 86.2 2.8 96.6 96.6 93.5 93.5 93.5
unconditional generation
Geo2Seq 82.5 n.a. 96.1 n.a. n.a. n.a. n.a.
QUETZAL 86.7 2.5 95.7 95.4 90.1 89.8 89.8
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1
NEAT (ours) 81.1 0.52 97.9 97.4 74.2 73.7 73.7
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.3 $\pm$ 0.5 $\pm$ 0.6 $\pm$ 0.6
prefix completion
QUETZAL 69.0 0.1 27.8 13.6 28.1 17.7 17.7
$\pm$ 1.8 $\pm$ 0.1 $\pm$ 6.8 $\pm$ 4.1 $\pm$ 4.7 $\pm$ 2.8 $\pm$ 2.8
NEAT (ours) 81.1 0.9 95.7 94.0 65.3 64.1 64.1
$\pm$ 1.1 $\pm$ 0.4 $\pm$ 2.0 $\pm$ 2.0 $\pm$ 3.2 $\pm$ 3.2 $\pm$ 3.2
@@
Figure 3: Randomly selected examples of molecules generated by NEAT trained on GEOM (white: hydrogen, gray: carbon, blue: nitrogen, red: oxygen, green: chlorine).
Figure 4: Top panel: Schematic workflow of the prefix completion task. Bottom panel: Randomly selected examples of complete molecules generated from three different prefixes: para-substituted benzene, 2-substituted furan, and 1,3-substituted indole.
Prefix completion
A common task in medicinal chemistry is to design molecules around a given core scaffold. We refer to this as prefix completion: given a target core, the model must produce valid completions consistent with that prefix. Existing approaches are poorly suited to this setting. Canonicalization-based methods such as Geo2Seq, cannot continue from arbitrary prefixes, since the prefix must coincide with the first $m$ atoms of the fully generated molecule after canonicalization. This information is unavailable prior to generation. In principle, permutation-invariant GNNs such as G-SchNet could support arbitrary prefixes, but in practice their generative quality is inferior and sampling is slow. Diffusion models such as EDM require special handling, since they assume joint denoising of all atoms. Xie et al. (2024) trained a diffusion model specifically for scaffold completion, but their approach relies on specialized architectural components and a tailored training procedure. Although Cheng et al. (2025) report that QUETZAL can perform prefix completion, they provide only qualitative results. Moreover, QUETZAL’s performance is sensitive to the atom ordering in the XYZ files used for training. In ablation studies, the authors showed that perturbing this order degrades performance, which is problematic for prefix completion.
Unlike prior autoregressive models, NEAT enforces atom-level permutation invariance and avoids sequential positional encodings, which is critical for effective prefix completion. To demonstrate this, we tasked NEAT and QUETZAL (trained on GEOM and run with the default settings of the best-performing model) with generating 1,000 molecules from 100 incomplete prefixes automatically extracted from the GEOM training set. These prefixes consist in common ring systems with different substitution patterns. SD files with the 3D atomic coordinates for all prefixes can be found in our code repository and 2D plots are provided in Appendix A.5. Prior to completion, each prefix was randomly rotated and translated for every generated sample. This procedure ensures that repeated generations from the same prefix do not share identical initial atomic coordinates, which could otherwise lead to similar outputs despite the stochasticity of the generation process. As shown in Table 2, NEAT significantly outperforms QUETZAL across all metrics. A summary of the prefix completion workflow as well as randomly selected examples of completed prefixes can be seen in Figure 4.
Runtime
Cheng et al. (2025) showed that combining a transformer with a diffusion loss substantially accelerates 3D molecular generation relative to diffusion models. This advantage extends to NEAT. Generating $10,000$ GEOM-like molecules takes $340$ s with NEAT versus $712$ s with QUETZAL. The runtimes were averaged over three runs on an NVIDIA GeForce RTX 4090 GPU with a batch size of 1000. To maximize comparability, both models use $N=60$ integration steps in this comparison.
@@
NEAT inherits standard limitations of autoregressive architectures. Error propagation might degrade sampling for larger molecules, which could be mitigated by a self-correction mechanism in future work. NEAT produces variable-sized molecules, but does not extrapolate far beyond the maximum atom count seen during training. For large molecules, pooling node embeddings may cause oversquashing and lose key information. Because bonds are not predicted explicitly, bond assignments still rely on predefined rules. Finally, despite random rotations during training, the lack of $SE(3)$ equivariance can yield orientation-dependent predictions.
@@ 5 Conclusion
We introduce a new method for training an autoregressive set transformer with bidirectional attention for 3D molecular generation using neighborhood guidance. Instead of sequential positional encodings, we rely on a subgraph sampling schema for model supervision. Treating the input as a permutation-invariant set removes the need for sequence canonicalization and makes the model well-suited for tasks like prefix completion. On QM9 and GEOM, our approach achieves competitive performance, while offering permutation-invariance with respect to atoms and fast sampling. We believe this provides a strong foundation for future work in conditional 3D molecular generation.
Impact Statement
This paper aims to advance the field of machine learning in general, and of de novo molecular design in particular. While we recognize the potential misuse of our approach to generate harmful compounds, we emphasize that there are numerous beneficial applications, such as for instance the development of new pharmaceuticals and safer agrochemicals. Importantly, the generation of molecules is only one step in a complex pipeline, requiring extensive validation, synthesis, and testing before any practical use, which mitigates immediate risks. Furthermore, the ethical deployment of such models should be guided by responsible practices, including rigorous oversight and adherence to safety standards. We do not foresee ethical concerns beyond those generally associated with generative models, but we stress the importance of responsible use to maximize societal benefit while minimizing risks.
@@
The source code and model weights are available at https://github.com/molinfo-vienna/NEAT.